OpenStudy (anonymous):

The magnitude of a star is modeled by M=6-(5/2)log(I/Io), where Io is the intensity of a just-visible star and I is the actual intensity of the star being measured. The dimmest stars of magnitude 6, and the brightest 1. Determine the ratio of light between stars of magnitude 1 and 3. (thanks!)

OpenStudy (anonymous):

You're not going to do any work yourself?

OpenStudy (anonymous):

Let the intensities be \(I_3\) and \(I_1\).\[ \frac{I_3}{I_1}=\frac{I_3}{I_0}\div \frac{I_1}{I_0} \]

OpenStudy (anonymous):

\[ \log\left(\frac{I_3}{I_1}\right)=\log\left(\frac{I_3}{I_0}\div \frac{I_1}{I_0} \right)=\log\left(\frac{I_3}{I_0}\right)-\log\left(\frac{I_1}{I_0} \right) \]

OpenStudy (anonymous):

\[ 3 = 6-\frac{5}{2}\log\left(\frac{I_3}{I_0}\right) \]And \[ 1 = 6-\frac{5}{2}\log\left(\frac{I_1}{I_0}\right) \]Subtract them both to get \[ 2 = \frac{5}{2}\log\left(\frac{I_3}{I_0}\right)-\frac{5}{2}\log\left(\frac{I_1}{I_0}\right) = \frac{5}{2}\left(\log\left(\frac{I_3}{I_0}\right)-\log\left(\frac{I_1}{I_0}\right) \right) \]

OpenStudy (anonymous):

Now try to do the rest on your own.

OpenStudy (anonymous):

Remember you are trying to find \(I_3/I_1\)

OpenStudy (wolf1728):

Each change in magnitude is a brightness difference of the 5th root of 100 or about 2.511556... So the brightness difference between magnitude 1 and 3 is equal to (2.511556...)² or 6.3079135

OpenStudy (anonymous):

Those 2 were the only ones I need help with, thanks so much for your time and help!